# Automatic Link Selection in Multi-Channel Multiple Access with Link Failures

### v1

@@ 9       Update the virtual queues
Below, we introduce the ROUND function, a technique adapted from the one introduced in [33] to approximate a nonnegative matrix by a matrix in the transport polytope. The technique can be readily extended to the Birkhoff polytope using the same algorithm.
@@
It can be shown that ROUND$(\boldsymbol{P})\in\mathcal{S}_{\varepsilon/s}^{\text{doub}}$ whenever $\boldsymbol{P}\in\mathcal{S}_{\varepsilon}$.
In Algorithm 1, we provide the algorithm for the task. In the following section, we focus on solving the intermediate problem (8), after which we move on to the analysis of the Algorithm. Problem (P2) motivates the definition of auxiliary variables $\boldsymbol{\gamma}(1),\boldsymbol{\gamma}(2),\dots\in[0,1]^{n}$ and a virtual queue $\boldsymbol{Q}(1),\boldsymbol{Q}(2),\dots$ to aid the algorithm. In particular, we will show that we can use our algorithm to get arbitrarily close to $\phi^{\text{opt}}$ in objective value.
III-A Solving Problem (8)
Finding $\boldsymbol{\gamma}(t+1)$: Notice that this problem can be solved due to the Assumption A2.
Finding $\boldsymbol{\tilde{P}}(t+1)$: We will only consider the case when $t$ is even. The case when $t$ is odd can be solved similarly. Notice that we can separately solve for each column of $\boldsymbol{\tilde{P}}(t+1)$. To solve for the $j$-th column of $\boldsymbol{\tilde{P}}(t+1)$, we define $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{s}$, where
[…]
where $D_{\text{KL}}(\cdot\lVert\cdot)$ in this case is the KL-divergence. It should be noted that (P3) has a classic structure that is solved in [11]. In particular, we define $\boldsymbol{z}$, where $z_{i}=y_{i}\exp(x_{i})$. First, assume that $\boldsymbol{z}$ is sorted in the increasing order. Then it can be shown that there exists $i\in[0:s-1]$ such that the vector $\boldsymbol{u}^{i}\in\mathbb{R}^{s}$ given by,
$u^{i}_{j}=\begin{cases}\varepsilon&\text{ if }j\leq i\\ \frac{z_{j}}{\sum_{l=i+1}^{s}z_{l}}(1-\varepsilon i)&\text{ if }j>i\\ \end{cases}$
We first establish the following two lemmas regarding the ROUND operation. A version of the following lemma is proved for the transport polytope in [33]. We adapt the idea for the Birkhoff polytope.
@@ III-A Solving Problem (8)
The input $\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}$, intermediate matrices $\boldsymbol{P}^{{}^{\prime}},\boldsymbol{P}^{{}^{\prime\prime}}$ and the output matrix $\boldsymbol{Q}$ of the ROUND function in Algorithm 1 satisfy
@@
    $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$.
  4. 4.
@@
See Appendix A. ∎
Next, we introduce the following lemma that bounds the difference $\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)$.
@@
Now, we introduce some useful preliminary lemmas. First, we introduce the push-back lemma regarding minimizing strongly convex functions (See, for example [51]).
Lemma 4 (Push-back)
Consider $\mathcal{Q}\in\{\mathcal{S}^{\text{doub}}_{\varepsilon},\mathcal{S}^{\text{row}}_{\varepsilon},\mathcal{S}^{\text{col}}_{\varepsilon},\mathcal{S}\}$, and $\varepsilon\in[0,1/s]$. Let $g:\mathbb{R}_{+}^{s\times s}\to\mathbb{R}$ be a convex function. Fix $\alpha>0$, and $\boldsymbol{Y}\in\mathcal{Q}$. Let
$\boldsymbol{X}^{*}\in\arg\min_{\boldsymbol{X}\in\mathcal{Q}}\left[g(\boldsymbol{X})+\alpha D(\boldsymbol{X}\lVert\boldsymbol{Y})\right],$
where $D(\cdot\lVert\cdot)$ is the divergence defined in (15). Then,
$g(\boldsymbol{X}^{*})+\alpha D(\boldsymbol{X}^{*}\lVert\boldsymbol{Y})\leq g(\boldsymbol{Z})+\alpha D(\boldsymbol{Z}\lVert\boldsymbol{Y})-\alpha D(\boldsymbol{Z}\lVert\boldsymbol{X}^{*}),$
for all $\boldsymbol{Z}\in\mathcal{Q}$.
We now establish the following two lemmas.
@@
We have for $\boldsymbol{X},\boldsymbol{Y}\in\mathcal{S}$,
where $D(\boldsymbol{X}\lVert\boldsymbol{Y})$ is the divergence defined in (15).
Proof:
Notice that
Next, we state a result on the properties of divergence D.
Lemma 6
Proof:
Notice that
$D(\boldsymbol{X}\lVert\boldsymbol{Y})$ $=\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{X_{i,j}}{Y_{i,j}}\right)\leq_{(a)}\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{1}{\varepsilon}\right)=s\log\left(\frac{1}{\varepsilon}\right),$
Before moving on to the main theorem, we establish a deterministic bound on the queue size $\lVert\boldsymbol{Q}(t)\rVert$ for our Algorithm (Algorithm 1).
Lemma 7
We have for all $t\in\mathbb{N}$ and $i\in[n]$, $Q_{i}(t)\leq BV+1$, where $B$ is the bound on the subgradients of $\phi$ defined in assumption A1 and $V$ is the utility parameter of Algorithm 1.
Proof:
We first prove the following claim.
Claim: If $Q_{k}(t)>BV$, for $k\in[n]$, then $\gamma_{k}(t+1)=0$.
Proof:
Assume the contrary $Q_{k}(t)>BV$, and $\gamma_{k}(t+1)>0$.
Consider $\boldsymbol{\tilde{\gamma}}$ such that
$\tilde{\gamma}_{i}=\begin{cases}0&\text{ if }i=k\\ \gamma_{i}(t+1)&\text{ otherwise.}\end{cases}$
Since, $Q_{k}(t)>BV$, we have
$[-V\phi^{{}^{\prime}}_{k}(\boldsymbol{\tilde{\gamma}})+Q_{k}(t)]>0,$ (21)
due to the bounded subgradient property (Assumption A1). Hence, notice that
$\sum_{i=1}^{n}$ $[-V\phi^{{}^{\prime}}_{i}(\boldsymbol{\tilde{\gamma}})+Q_{i}(t)](\gamma_{i}(t+1)-\tilde{\gamma}_{i})=\gamma_{k}(t+1)[-V\phi^{{}^{\prime}}_{k}(\boldsymbol{\tilde{\gamma}})+Q_{k}(t)]>0,$ (22)
where the last inequality follows due to $\gamma_{k}(t+1)>0$ from assumption and (21).
The subgradient inequality for the convex function $f:[0,1]^{n}\to\mathbb{R}$ given by, $f(\boldsymbol{\gamma})=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}$ gives,
$\sum_{i=1}^{n}[-V\phi^{{}^{\prime}}_{i}(\boldsymbol{\tilde{\gamma}})+Q_{i}(t)](\gamma_{i}(t+1)-\tilde{\gamma}_{i})\leq f(\boldsymbol{\gamma}(t+1))-f(\boldsymbol{\tilde{\gamma}})\leq 0,$
where the last inequality follows from the optimality of $\boldsymbol{\gamma}(t+1)$ for (P3). This is a contradiction with (22). ∎
Now, we move on to the proof of Lemma 7. Fix $i\in[n]$. We use induction to prove that $Q_{i}(t)\leq BV+1$ for all $t\in\mathbb{N}$. Notice that the result trivially follows for $t=1$. Assume $Q_{i}(t)\leq BV+1$. We establish $Q_{i}(t+1)\leq BV+1$. We consider two cases.
Case 1: $Q_{i}(t)>BV$: Notice that from the previous claim, we should have $\gamma_{i}(t+1)=0$. Hence, we have from the queuing equation (16),
$Q_{i}(t+1)$ $\leq Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\leq_{(a)}Q_{i}(t)\leq BV+1,$
where (a) follows since $\gamma_{i}(t+1)=0$, and $X_{i}(t)\geq 0$, and the last inequality follows from induction hypothesis.
Case 2: $Q_{i}(t)\leq BV$: We have from the queuing equation (16),
$Q_{i}(t+1)$ $\leq Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\leq_{(a)}Q_{i}(t)+1\leq BV+1,$
where (a) follows since $\gamma_{i}(t+1)\leq 1$, and $X_{i}(t)\geq 0$, and the last inequality follows from the description of the case.
Hence, we are done. ∎
Define the drift $\Delta(t)$ as
$\Delta(t)=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
and the history $\mathcal{H}(t)$ as the sigma algebra generated by
[…]
Notice that $\boldsymbol{\tilde{P}}(\tau),\boldsymbol{P}(\tau),\boldsymbol{\gamma}(\tau),\boldsymbol{Q}(\tau)$ for $\tau\in[t]$ are $\mathcal{H}(t)$-measurable.
Now, we have the following bound on $\Delta(t)$.
Lemma 8
We have for all $t\in\mathbb{N}$,
Proof:
Notice that from the queuing equation (16), we have for all $i\in[n]$,
$Q^{2}_{i}(t+1)$ $\leq\left(Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\right)^{2}$
$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right],$
Taking the expectations conditioned $\mathcal{H}(t)$:
Taking expectations, and performing simple algebraic manipulations, we have the result. ∎
Lemma 9
We have
$-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\}$ $\leq\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}$
$\ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\},$
@@
$=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)]\right\}+\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t)-P_{i,j}(t)]\right\}$
$\ \ \ \ +\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}.$ (24)
@@
$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{12s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t)-\boldsymbol{P}(t)\rVert^{2}_{1}\right\}$
$\leq_{(a)}3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{3s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert^{2}_{1}\right\}$
$\leq_{(b)}3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{2}{3\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\},$
where (a) follows from Lemma 3, and (b) follows from Lemma 5. Next, notice that
$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}$ $=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\frac{S_{i,j}(t)}{P_{i,j}(t)}P_{i,j}(t)Y_{i,j}(t)\right\}$
$=\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\},$
@@
where (a) follows since $P_{i,j}(t)\geq\varepsilon/s$ from Lemma 3, (b) follows since $q_{i,j}\leq 1$, and the last inequality follows due to Lemma 7. Combining with (9), we are done. ∎
@@
Adding $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 8, we have
@@
where (a) follows by Lemma 9. Now notice that from the optimality of $\boldsymbol{\tilde{P}}(t+1),\boldsymbol{\gamma}(t+1)$ in (8) (see Algorithm 1) with Lemma 4, we have for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\varepsilon}$,
@@
$\ \ \ \ =-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}P_{i,j}\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}$
@@
where for the last inequality, we have used Lemma 6. Rearranging the above, and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$ from Lemma 7, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done. ∎
Lemma 11
We have for any $T,T_{0}\in\mathbb{N}$,
Proof:
Notice that
$\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}+1}^{T+T_{0}}\boldsymbol{\gamma}(t)\right)\right)$
$=\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)\right)\right)$
$\leq_{(a)}\phi\left(\frac{1}{T}\Bigg{(}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}[\boldsymbol{Q}(t+1)-\boldsymbol{Q}(t)]\right)\Bigg{)}$
where (a) follows from the entrywise nondecreasing property of $\phi$, and the queuing equation (16), for (b) we use $\boldsymbol{Q}(T_{0})\geq 0$ and the entrywise nondecreasing property of $\phi$, and (c) follows since $\phi$ is $\sqrt{n}B$-Lipschitz continuous from Assumption A1.  [^2] Assumption A1 ensures that $\phi$ can be extended to an entrywise nondecreasing function that is $\sqrt{nB}$-Lipschitz continuous over the domain $[0,\infty)^{n}$.
Now, we are ready to establish the performance bound of the algorithm.
Theorem 1
For any parameters $T,T_{0}\in\mathbb{N}$, $\varepsilon\in(0,1/s)$, $\eta,V>0$, Algorithm 1 yields
$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)$ $\leq\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}$
$\ \ \ \ +\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}.$
In particular
  1. 1.
    For any $\epsilon>0$, choosing $V=\Theta(1/\epsilon)$, $\varepsilon=\Theta(\epsilon)<1/s$, $\eta=\Theta(\epsilon^{3})$, we have
    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)=O(\epsilon).$
  2. 2.
    In the finite time horizon setting with $T,T_{0}\in\mathbb{N}$, using $\eta=\Theta(1/T)$, $\varepsilon=\Theta(1/T^{1/3})<1/s$, and $V=\Theta(T^{1/3})$, we have
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\frac{\log(T)}{T^{1/3}}\right).$
@@
where (a) follows since $\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\geq\gamma_{i}^{*}$ for all $i\in[n]$ from the feasibility of $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ for (P2), and Lemma 7. Now, we divide both sides of the above inequality by $VT$ and use the Jensen’s inequality to obtain

### v2

@@ V-A ROUND function
We introduce ROUND, a technique adapted from the one introduced in [5] to approximate a nonnegative matrix by a matrix in the transport polytope. The technique can be readily extended to the Birkhoff polytope.
@@
We establish the relationships between the input and the output matrices of the ROUND operation. A version of the following lemma is proved for the transport polytope in [5]. We adapt the idea for the Birkhoff polytope.
@@
The input $\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}$ and the output matrix $\boldsymbol{Q}$ of the ROUND function satisfy
@@
See Appendix C
V-B Analysis of the Algorithm
The rest of the section focuses on proving Theorem 2. Similar to Section IV, we consider the interval $T_{0}:T_{0}+T-1]$, where the success probabilities $q_{i,j}$ remain fixed. We first notice that Lemma 6, Corollary 1.1, and Lemma 8 hold in the same form for Algorithm 2. These results are used in the proof of Theorem 2. The intermediate problem (55) admits an explicit solution for $\boldsymbol{\tilde{P}}(t+1)$, as discussed in Appendix D.
First, we introduce the following lemma that bounds the difference $\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)$.
@@ V-B Analysis of the Algorithm
Now, we first introduce the following lemma that bounds the error introduced by replacing the probabilities $q_{i,j}$ with $\hat{S}_{i,j}(t)$.
@@
We have for any $t\in[T_{0},\,T_{0}+T-1]$,
- $\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\}$
$\leq\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}-\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\},$
@@
$=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)]\right\}$
$\ \ \ \ +\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t)-P_{i,j}(t)]\right\}+\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}.$ (60)
@@
$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t)-\boldsymbol{P}(t)\rVert^{2}_{1}\right\}}{12s\eta}$
$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert^{2}_{1}\right\}}{3s\eta}$
$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{2\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}}{3\eta},$
where the second inequality follows from Lemma 13, and the last inequality follows from Lemma 3. Next, notice that
$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}$
$=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\frac{S_{i,j}(t)}{P_{i,j}(t)}P_{i,j}(t)Y_{i,j}(t)\right\}=\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\},$
@@
where (a) follows since $P_{i,j}(t)\geq\varepsilon/s$ from Lemma 13, (b) follows since $q_{i,j}\leq 1$, and the last inequality follows due to Corollary 1.1. Combining with (14), we are done. ∎
@@
Adding $V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 8, we have
@@
where (a) follows by Lemma 14. Now notice that from the optimality of $\boldsymbol{\tilde{P}}(t+1),\boldsymbol{\gamma}(t+1)$ in (55) (see Algorithm 2) with Lemma 2-1, we have for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\varepsilon}$,
@@
$=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}-\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}$
@@
where for the last inequality, we have used Lemma 4. Rearranging the above, and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$ from Corollary 1.1, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done. ∎
Now, we are ready to establish Theorem 2
@@
where (a) follows since $\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\geq\gamma_{i}^{*}$ for all $i\in[n]$ from the feasibility of $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ for (P2), and Corollary 1.1. Now, we divide both sides of the above inequality by $VT$ and use the Jensen’s inequality to obtain